# Biomechanics-Aware Trajectory Optimization for Navigation during Robotic Physiotherapy

### v1

@@ II-E Robot control and model-based gravity compensation
Equations (11)-(12) ensure that the robot’s end-effector pose is consistent with the desired trajectory for the human. Given the optimal shoulder angles, the EE pose follows from the human elbow position and orientation, with the end-effector orientated so that the upper arm always points to the center of the human shoulder. The end-effector’s reference poses are computed coherently with the frequency of the NLP solution, meaning that a new pose is commanded to the robot every $T_{f}/N$ seconds. A Cartesian impedance controller, running at $200$ Hz, tracks the target end-effector poses. The stiffness and damping parameters of the controller were tuned experimentally.
However, commanding the Cartesian end-effector pose through (12) produces non-negligible tracking errors along the vertical axis. This issue is primarily due to the arm’s weight, since early rehabilitation is typically administered with the human subject not yet able to support the weight of their own arm [44]. Impractically high stiffness levels, far exceeding what is safe for the subject, would be required to significantly improve the tracking. To remedy this, we implemented a gravity-compensation strategy that integrates the generalized torques $\boldsymbol{u}^{*}$ into the end-effector reference. More specifically, the vertical reference for the end-effector is adjusted by including the optimal value $u^{*,SE}$ of the vertical movements along the $SE$ DoF, which account for arm weight and inertia through (3). This additional torque is transformed into the corresponding vertical force to be applied at the human elbow. Given the desired stiffness of the impedance controller along the vertical axis ($K_{z}$), such a force can be realized at the end-effector by modulating the distance of the attractor from the current position:
${}^{\text{base}}\bar{p}_{\text{EE}}^{z}={}^{\text{base}}\hat{p}_{\text{EE}}^{z}+\frac{u^{*,\text{SE}}}{K_{z}L_{\text{hum}}\mathrm{sin}(\hat{\text{SE}})}$ (13)
This compensation adjusts the robot’s vertical reference, achieving the desired height set-point and fully supporting the human arm without high stiffness. The vertical force produced by the robot counteracts gravity on the arm at the desired elbow height, naturally descending from the personalized musculoskeletal model used in the trajectory optimization. Additionally, the formulation in (13) is robust to model inaccuracies: if the arm’s actual weight differs from expectations, the impedance controller will stabilize at an equilibrium point. This is advantageous with respect to feedforward methods (i.e. [45]) that actively compensate for gravity, requiring accurate estimates of the subject’s mass and its distribution.
II-F Experimental setup and design
Our experimental setup included a KUKA LBR iiwa 7 robotic arm equipped with a custom-built brace that interfaces the human with the robot end-effector. The brace was shaped with a thermoplastic material to cradle the human elbow (see Fig. 3), and attached to the robot end-effector via screws, fixing the transformation between the robot EE frame and the human elbow frame. Through this elbow brace interface, the robot could impose a desired movement on the human and support the weight of their arm, while simultaneously estimating the human state. Computations were performed on a Dell Latitude 7420 laptop with an i7-1185G7 processor, interfaced with a dedicated Dell workstation with an Xeon W-2123 processor running the impedance controller. In this way, we ensured that the robot low-level controllers would receive commands at the required 200 Hz frequencies, regardless of potential delays in the high-level planning.  [^2] Prior to paper acceptance, our software implementation is available by contacting the authors.
@@ II-F Experimental setup and design
A healthy human subject acted as a patient in our lab-based experiments. The brace interface was strapped onto the subject’s arm, and they were instructed to maintain a constant pose for their torso with respect to the robot base. This satisfied the assumption of a stationary glenohumeral joint center employed in the model’s state estimation. Our experimental protocol was approved by the TU Delft Human Research Ethics Committee (HREC).
To demonstrate the capabilities of our approach, we considered two scenarios: the early stages of the therapy, where the robot performs full guidance while the human subject is passively led through the movements, and the later stages of physical therapy, where the human subject gradually regains strength and actively participates in the movements. We describe each scenario in more detail below.
II-F1 Passive human subject
when the subject passively follows the movement, their rotator cuff tendon strain is purely position-dependent. To guarantee low strains across all of the rotator cuff, tendon behavior was monitored through a strain map aggregating information regarding the maximum strain across all of the tendons. In other words, the strain value at a given model state $\boldsymbol{x}$ represents the maximum strain observed across the supraspinatus, infraspinatus, subscapularis, and teres minor tendons. We locked the $AR$ DoF so that the complete rehabilitation movement towards the target final state $\boldsymbol{x}_{d}$ can be optimized and then executed in an open loop. We fixed the time horizon for the optimization to $T_{f}=5$ seconds and divided it into $N=50$ equal intervals. We determined the cost function weights for the real human-robot experiment via extensive simulations, to provide the potential user of our approach with general guidelines and understanding of parameter selection with respect to the desired behavior.
II-F2 Active human subject
due to active human response during therapy, unforeseen modifications to the strain maps are to be taken into account. In particular, we are interested in how specific muscle activations can be accounted for in our pipeline. Without loss of generality, we present here therapy targeting the supraspinatus tendon specifically, selected as it is one of the most commonly injured within the rotator cuff [46]. As the supraspinatus muscle activation $\alpha_{\text{ss}}$ modifies tendon strain, our trajectory optimization must be reactive to adapt the optimal trajectory in real-time over the evolving strain map. As such, we considered a shorter planning horizon that allows us to compute and update the optimal trajectory in real-time ($T_{f}=1$ second, $N=10$ intervals), optimizing it over a receding horizon. First, we demonstrated in simulation the reactiveness of our trajectory optimization to different strain map variations, and then moved to closed-loop execution on the real robot. In this experiment, we left the AR DoF free, as our real-time replanning is able to handle generic variations to the 2D strain maps.
In a previous study, we demonstrated that surface electromyography (EMG) does not accurately reflect rotator cuff activity and strain [22]. Therefore, instead of using EMG to measure muscle activations, we emulated them and assessed their effect on tendon strain using OpenSim. During our experiments, the emulated activation was kept constant, presented sudden steps, or followed ramp profiles. These cases were considered as they cover conceptually both fast and slow variations in the activation of the subject’s muscles during therapy. In the future, realistic activation estimates could be achieved with a real-time algorithm like our rapid muscle redundancy (RMR) solver [22]. Given the time-consuming setup and potential reliability issues of EMG, model-based or hybrid muscle activity estimation would be more practical for real physiotherapy applications.
III Results
@@
III-A Effect of OCP parameters on trajectories (passive subject)
@@
Figure 5: Human trajectories produced for passive rehabilitation were visualized on the corresponding strain map (left). The blue line shows the actual trajectory, in human coordinates, along which the robot moved our subject. The paths connect target human poses given as goals to 3 instances of our trajectory optimization problem minimizing the maximum strain that the rotator cuff tendons encounter. The corresponding pose of the test subject is shown at the different target positions in the motion (right).
@@
We first ran BATON in simulation to examine the effects of optimization parameters and to determine the best relative cost function weights for synthesizing rehabilitation movements for passively assisted patients. We explored various permutations of the cost function weights ($w_{\sigma}$, $w_{\text{acc}}$, $w_{\tau}$), with Fig. 4 showing some of the representative scenarios. We fixed $w_{\text{acc}}=10$ to guarantee low accelerations on the human body and varied the other two parameters to achieve rather extreme cases. If no requirement of minimizing the strain is specified ($w_{\sigma}=0$), our optimization would drive the human arm towards the target goal pose along the shortest path (dashed blue line). When higher weights $w_{\sigma}$ are used, the trajectory makes a wider excursion in the low-strain regions (solid and dotted blue lines). While this guarantees safer therapy, we observed that increasing this term gradually produces wider trajectories while the quantitative decrease in strain becomes negligible. In other words, wider excursions do minimize the cost function value, but they lead toinsignificant reductions in strain. Given this trade-off, we fixed $w_{\sigma}=1$ for later experiments.
@@
With a passive subject, the strain map was fixed and did not vary with time. Thus, we could directly optimize the entire trajectory and execute it in an open-loop manner. In these conditions, the terminal condition in II-B2 is applied also to the final human pose, so the optimized trajectory reaches it even when $w_{\tau}=0$. This effectively simplifies the tuning of our cost function to only two parameters in this case. By selecting $w_{\text{acc}}=10$, $w_{\sigma}=1$ the resulting trajectory traverses safe low-strain regions ($\sigma\leq 2\%)$. These weights were used for the robot experiments in Sec. III-B below.
@@
However, relying only on the terminal constraint to drive the human body to the target pose is not feasible when variations in the strain map can happen. Indeed, a trajectory that is safe on the initial map can end up crossing high-strain regions unexpectedly on the updated strain map. This motivates setting $w_{\tau}\neq 0$, and investigating the effect of normalizing the distance to the target $L_{\tau}$ on the optimal trajectory (Fig. 4, green lines). When $L_{\tau}$ was normalized by the initial distance from $\boldsymbol{x}_{\tau}$, lower-strain trajectories were produced, prioritizing strain minimization over approaching the goal earlier. This normalization ensured that the goal pose always provides a movement direction, while preventing the contribution to the cost function $L$ from vanishing when the goal is closer or dominating when it is farther.
All the trajectories were simulated on the KUKA robot in Gazebo, to check for the robot’s ability to reproduce the desired movements.
III-B Quality of the executed paths (passive subject)
@@
To demonstrate the performance of BATON in delivering therapy motion to a test subject by a real robot, we selected a few target poses around the shoulder state space to which the robot had to drive the subject. Figure 5 shows the experimental setup and the target poses on the strain map, together with the optimal trajectories that were delivered to our test subject by the robot. The robot started at the state marked by the blue dot in the low-strain zone. First, it moved the patient to the state marked by the orange dot and then toward the state indicated by the green dot to increase the level of strain and increase the range of motion. The resulting trajectory avoided higher strain regions and lower overall strain was achieved in the movement (Fig. 6, top row), permitting the exploration of a large range of motion with minimum risk of re-injury. Finally, the robot moved the shoulder state back into the low-strain zone denoted by the pink dot.
While human safety is more easily analyzed in the subject-centric coordinates on the strain map, the impedance controller tracks the desired optimal trajectory expressed in its base frame. The Cartesian reference positions for the robot end-effector are presented in Fig. 6 (last two rows), together with the actual end-effector poses across the duration of the experiment. In particular, the bottom row of the figure presents the tracking performance along the robot’s vertical axis. As expected, the vertical reference $z_{\text{}ref}$ is not reached, but the position mismatch $z_{\text{ref}}-z_{\text{curr}}$ allows the Cartesian impedance controller to provide the correct vertical force required to track the actual optimal end-effector height $z_{\text{opt}}$, which is obtained from our computed offset in (12).
Figure 6: Evolution of the strain level $\sigma$ experienced by the human during passive therapy (above) and Cartesian coordinates of the robotic end-effector ${}^{\text{base}}\boldsymbol{p}_{\text{EE}}$ during the experiment (below). The top graph compares $\sigma$ during the experiment against the strain along the shortest paths between the beginning and end of each trajectory. The bottom graph highlights the effect of using the personalized musculoskeletal model to inform the gravity compensation of the human arm: the reference given to the controller (solid black) makes it so that the current position (dashed) tracks the real optimal one (solid blue).
III-C Online adaptations (active subject)
@@
Figure 7: The simulated effects of supraspinatus muscle activation ($\alpha_{\text{ss}}$) on the optimal trajectory towards the same goal pose, when $\alpha_{\text{ss}}$ follows different activation profiles (green: decreasing ramp, white: step) or remains constant (blue). Large activation steps are used for ease of visualization, and all of the resulting trajectories are projected on the strain map corresponding to the highest $\alpha_{\text{ss}}$ for comparison to the right. Our musculoskeletal trajectory optimization is capable of accounting for sudden variations in the strain level, as a consequence, e.g., of varying muscle activation.
III-C1 Simulations
@@
We investigated the responsiveness and adaptability of our control scheme when the strain maps change as a consequence of varying muscle activations. Muscle activation has a direct effect on the optimal path between two poses, since it affects the strain on the corresponding tendon, through (2). We studied this functionality in simulation to explore how various muscle activation patterns affect our rehabilitation trajectory, and to test our optimization with more drastic and challenging strain map variations. The initial model state was set at $SE_{0}=PE_{0}=60^{\circ}$, while the target position to reach was set to $\bar{PE}=45^{\circ}$, $\bar{SE}=95^{\circ}$, requiring movement along multiple human DoFs. Owing to the shorter time horizon over which we ran our trajectory optimization, BATON was able to achieve a trajectory re-planning with an update rate of roughly 10 Hz. This permits us to generate new trajectories that can account for strain map variations, allowing us to control human movement in a closed-loop manner. We varied the (artificial) muscle activation level for the supraspinatus muscle to assess how the strain modifications affect the resulting trajectories.
@@
Fig. 7 demonstrates the effect of muscle activation variations: a sudden step increase in activation (white trajectory) and a gradual decrease in activation (green trajectory). These two cases are shown alongside the two boundary cases where activation is kept constantly high or low (resulting in the blue trajectories in the figure). On the right side of Fig. 7, we visualize all of the trajectories jointly on a single strain map, representing the strain landscape corresponding to the highest activation considered. Re-projecting all of the trajectories on the same plane allows one to see the different human trajectories produced by the gradual and sudden variations of the activation level, bounded by the two cases in which the activation was kept constant.
@@
Figure 8: Real-time trajectory optimization allows to account for sudden variations in the strain level. We show here how the optimal trajectory changes as a consequence of rapid variations in a (fabricated) strain map. The trajectory optimization starts on a strain map which is completely safe, generating a motion directly towards the goal (left). Sudden modifications to the strain cause the reference to avoid higher-strain zones (center and right), generating a trajectory that deviates from the shortest path adapting to the new scenarios, and prioritizing safety.
@@
To understand the effects of more sudden and large changes in activation, for example, due to a strong reflex response, we tested cases in which BATON is presented with larger, more localized, and sudden variations in the strain maps. In this case, the initial and target human poses are inverted, representing the previous movement that evolved in the opposite direction (see Fig. 8). First, the strain map on which to plan has low strains, so the optimal trajectory points directly to the goal (Fig. 8, left). We then introduced a sudden large local variation in the landscape, so that higher strain appears along the direction of movement. BATON reacted to the variation by re-planing a new trajectory that keeps evolving towards the lower-strain region, deviating from the older plan proposed at the previous time step (white dots in the figure). Similarly, we introduced a second sudden large local variation in the strain map, observing a new trajectory adaptation that allows the simulated movement to remain safe.
III-C2 Real system
having explored and compared the real-time adaptation of the reference optimal rehabilitation trajectories in simulation, we proceeded to demonstrate this in a real robotic-assisted experiment. The starting and the final human pose were the same as the first simulation experiment (respectively, $SE_{0}=PE_{0}=60^{\circ}$, and $\bar{PE}=45^{\circ}$, $\bar{SE}=95^{\circ}$). As explained above, the muscle activation $\alpha_{\text{ss}}$ was not directly measured by electromyography, but was emulated for the given task. The subject was guided through the same movement multiple times, during which the strain map varied to mimic interactive subject response due to time-dependent muscle activation. We demonstrate the real-time generation and tracking of the two extreme trajectories for $\alpha_{\text{ss}}=0$ and $\alpha_{\text{ss}}=0.25$ in Fig. 9, as simulation showed that the other cases would fall between these two. It can be noted that the algorithm can handle different levels of muscle activations by generating different rehabilitation movements, that are tracked in real-time by the Cartesian impedance controller. In particular, the two conditions lead to a difference of more than $10^{\circ}$ along the plane-of-elevation axis. This difference was a result of adaptations to the human movements to traverse safer areas defined by the different muscle activity values.
III-D Computational performance (active and passive subject)

### v2

@@ IV-A Robot control
Equations (14)-(15) ensure that the robot’s end-effector pose is consistent with the desired trajectory for the human. Given the optimal shoulder angles, the EE pose follows from the human elbow position and orientation, and is oriented such that the upper arm always points to the center of the human shoulder.
To track target end-effector poses, we used an impedance controller that governs the interaction force/torque in Cartesian space as:
$\boldsymbol{f}=\boldsymbol{K}(\bar{\boldsymbol{X}}_{\text{EE}}-\hat{\boldsymbol{X}}_{\text{EE}})-\boldsymbol{D}\hat{\dot{\boldsymbol{X}}}_{\text{EE}},$ (16)
where $\boldsymbol{f}$ is the interaction force/torque between the robot and its patient, $\bar{\boldsymbol{X}}_{\text{EE}}\in\mathbb{R}^{6}$ is the commanded EE reference retrieved from ${}^{\text{base}}\bar{\boldsymbol{R}}_{\text{EE}}$ and ${}^{\text{base}}\bar{\boldsymbol{p}}_{\text{EE}}$, and $\hat{\boldsymbol{X}}_{\text{EE}}\in\mathbb{R}^{6}$ is the measured actual pose of the robot, respectively. $\boldsymbol{K}\in\mathbb{R}^{6\times 6}$ and $\boldsymbol{D}\in\mathbb{R}^{6\times 6}$ are the Cartesian stiffness and damping matrix, respectively. The impedance controller operates at $200$ Hz.
In addition to the Kuka’s built-in gravity compensation, we include further compensation to support the mass of the subject’s arm. Specifically, we use the optimal generalized torques from the OCP in (7) to adjust the vertical reference for the robot’s EE, increasing it proportionally to the torques required to produce the optimal movement. This adjustment, further detailed in the Appendix, allows our system to keep the actual vertical position of the human arm close to the prescribed one, leveraging pure impedance control without the need to feed-forward assistive torques.
IV-B Experimental design for evaluation of BATON
Our experimental setup consisted of a KUKA LBR iiwa 7 robotic arm instrumented with a Bota SensONE force-torque (F/T) sensor (Bota Systems AG, Zurich, CH) and a custom thermoplastic elbow brace for interfacing the human subject with the robot. The brace was molded to cradle the subject’s elbow (Fig. 3) and mounted directly onto the F/T sensor at the robot’s end-effector, fixing the transformation between the robot EE frame and the human elbow frame. This configuration allowed the robot to impose motion on the human arm, support its weight, and simultaneously estimate the subject’s biomechanical state.
@@ IV-B Experimental design for evaluation of BATON
A healthy adult subject participated in our lab-based experiments, approved by the Human Research Ethics Committee at Delft University of Technology. The brace was strapped to the participant’s arm, and they were instructed to maintain a fixed torso posture relative to the robot base, satisfying the assumption of a stationary glenohumeral joint center required by our state estimation model.
To evaluate BATON and highlight the importance of modeling human arm dynamics in trajectory planning, we examined two rehabilitation scenarios: (1) the early stages of therapy, during which the robot provides full guidance while the subject remains passive; and (2) the later stages, where the subject progressively regains strength and actively participates in the motion. The first scenario also served to benchmark BATON against the state-of-the-art (SoTA) method proposed by Prendergast et al. [38], which generates strain-aware shoulder rehabilitation movements based on tendon behavior. Their method uses a modified A* algorithm to traverse discrete passive strain maps while minimizing cumulative strain, returning a kinematic sequence of optimal joint angles $\boldsymbol{\mathrm{q}}^{\text{SoTA},*}$. We implemented this planner in our framework and compared its performance with BATON’s in guiding a passive subject. The second scenario, involving an active subject, falls outside the applicability of the SoTA method, which assumes constant muscle activations and thus cannot accommodate unpredictable activation changes.
IV-B1 Passive human subject
In this scenario, the subject remained relaxed, allowing the robot to lead the movement. Muscle activity was negligible, and tendon strain was primarily determined by the shoulder pose. To ensure uniformly low strain across the rotator cuff, we monitored a strain map that aggregated the maximum strain values over the supraspinatus, infraspinatus, subscapularis, and teres minor tendons. The strain value at a given model state $\boldsymbol{x}$ was defined as the maximum of these individual tendon strains. To maintain consistency on the strain map, the axial rotation (AR) degree of freedom (DoF) was locked, so that the complete movement toward the target arm pose $\boldsymbol{q}_{\tau}$ evolved along a single map. Negligible muscle activation and constant axial rotation imply that trajectories can be optimized and then executed in an open loop. Multiple target poses $\mathbf{q}_{\tau}$ were defined and used as inputs to BATON to explore a broader range of motion at the glenohumeral joint. Each optimization instance used a time horizon of $T_{f}=5$ seconds, discretized into $N=50$ intervals. Cost function weights were selected based on extensive simulations to provide parameter tuning guidelines.
All candidate trajectories were first tested in the Gazebo simulator to verify the robot’s ability to execute them accurately. We then compared BATON-generated trajectories to those from the SoTA method, each executed on the physical robot. The SoTA trajectory $*\mathbf{q}_{\text{SoTA,*}}$ was upsampled to match the execution frequency of BATON’s output and fed to the controller via Eqs. (14)–(15). Performance was evaluated in terms of strain minimization, trajectory smoothness, and human–robot contact forces. Both methods started from the same human shoulder pose $\boldsymbol{q}(0)$ and navigated between the same target poses.
IV-B2 Active human subject
Due to human volitional actions and/or reflexive reactions during therapy, the level of muscle activity and forces can change quickly and significantly, impacting tendon strain estimates. BATON needed to adapt its trajectory planning in real time to maintain safety under these dynamic conditions. We shortened the planning horizon to $T_{f}=1$ second, with $N=10$ intervals, allowing for continuous replanning using a receding horizon scheme.
We first used simulations to investigate the reactiveness of BATON to artificial muscle activation changes when treating a specific tendon. Without loss of generality, the infraspinatus was considered. The initial human pose was set at $SE_{0}=PE_{0}=60^{\circ}$, while the target position to $\bar{PE}=45^{\circ}$, $\bar{SE}=95^{\circ}$, requiring movement along multiple human DoFs. Simulation also allowed us to consider drastic variations on fabricated maps to trigger more extreme trajectory adaptations as a consequence, for example, of a strong reflex response. In this case, the previous initial and target positions were inverted, representing the previous movement that evolved in the opposite direction.
Then, we moved to a closed-loop execution of BATON on the real robot, employing realistic infraspinatus activity estimation $\hat{\alpha}_{\text{is}}$ through the RMR solver (Sec. III-A).
In this scenario, we analyzed the effect that physical human-robot interaction and volitional human behavior have on online therapy adjustments. BATON guided our subject between the same initial and final poses ($SE_{0}=PE_{0}=60^{\circ}$, $\bar{PE}=45^{\circ}$, $\bar{SE}=95^{\circ}$) under two conditions: (i) the subject complied with the robot, and (ii) the subject actively engaged the infraspinatus by actively externally rotating their arm. This rotation, while not substantially disrupting the robot-guided trajectory, was resisted by the Cartesian controller, which maintained the initial value of the $AR$ DoF. The resulting interaction force elicited infraspinatus activation, consistent with its primary anatomical function [51].
Model-based computations were run on a Dell Latitude 7420 laptop with an i7-1185G7 CPU, interfaced with a workstation running a Xeon W-2123 and executing the robot’s impedance controller. This ensured low-level control at 200 Hz, decoupled from potential delays in high-level planning. The Robot Operating System managed communications between the various software modules.
V Results
We performed a thorough evaluation of the proposed methods. Regarding the planning aim (A1), we analyzed the effect of optimization parameters on the planned trajectory (Sec. V-A) and compared BATON to the SoTA method in terms of the quality of the executed paths (Sec. V-B). Regarding adapting to changing muscle activity (A2), we demonstrated BATON’s ability to re-plan online based on changing strain maps, first in simulation and then thanks to real-time muscle activity estimates (Sec. V-C). We concluded by conducting an analysis of real-time computational performance (Sec. V-D).
@@ V Results
V-A Effect of OCP parameters on trajectories (passive subject)
@@
Figure 5: Top left: human trajectories produced by our approach for passive rehabilitation (solid blue), visualized on the corresponding strain map together with the result of the planner in [38] (dashed grey). The two lines are the actual trajectory, in human coordinates, along which the robot moved our subject. The paths connect target human poses given as goals to 3 instances of our trajectory optimization problem minimizing the maximum strain that the rotator cuff tendons encounter, and 3 instances of the planner in [38]. Top right: goal poses of the test subject during the experiment. The graphs at the bottom display the magnitude of the acceleration along the two trajectories (left) and the interaction force magnitude between the robot and subject during the motion (right). Note that the data corresponding to re-planning periods is not shown for conciseness.
@@
We first ran BATON in simulation to examine the effects of optimization parameters and to determine the best relative cost function weights for synthesizing rehabilitation movements for passively assisted patients. We explored various permutations of the cost function weights $w_{\sigma}$, $w_{\text{acc}}$, and $w_{\tau}$, modulating the importance that tendon strain, human accelerations, and movement towards the goal pose have in the overall cost function. Fig. 4 shows some of the representative scenarios. We fixed $w_{\text{acc}}=10$ and varied the other two parameters to achieve rather extreme cases. If minimizing the strain is ignored ($w_{\sigma}=0$), the planner would choose the shortest path (dashed blue line). As higher weights $w_{\sigma}$ are used, the trajectory makes a wider excursion in the low-strain regions (solid and dotted blue lines). While this guarantees safer therapy, we observed that increasing this term gradually produces wider trajectories while the quantitative decrease in strain becomes negligible. In other words, wider excursions do minimize the cost function value, but they lead to insignificant reductions in strain. Given this trade-off, we fixed $w_{\sigma}=1$ for later experiments.
@@
With a passive subject, the strain map was fixed and did not vary with time. Thus, we could directly optimize the entire trajectory and execute it in an open-loop manner. Here, both final human pose and velocities are constrained (see Sec. II-B2), so the target pose is reached even when $w_{\tau}=0$. This terminal constraint effectively simplifies the tuning of our cost function to only two parameters in this case. By selecting $w_{\text{acc}}=10$, $w_{\sigma}=1$ the resulting trajectory traverses safe low-strain regions ($\sigma\leq 2\%)$. These weights were used for the robot experiments in Sec. V-B below.
@@
However, enforcing a terminal constraint on the final human pose is undesirable with shorter planning horizons, which are needed to react to changes in the strain map as muscle activations vary (active subject scenario, Sec. V-C). A trajectory deemed safe on the initial map may later intersect high-strain regions as the map is updated. This motivates setting $w_{\tau}\neq 0$, and investigating the effect of normalizing the distance to the target $L_{\tau}$ on the optimal trajectory (Fig. 4, green lines). When $L_{\tau}$ was normalized by the initial distance from $\boldsymbol{\mathrm{q}}_{\tau}$, lower-strain trajectories were produced, prioritizing strain minimization over approaching the goal earlier. This normalization ensures that the goal pose provides a movement direction, and prevents this information from vanishing when the goal is closer or dominating the cost when $\boldsymbol{\mathrm{q}}_{\tau}$ is farther.
V-B Quality of executed path w.r.t. SoTA
@@
Resulting trajectories from BATON were compared to the ones obtained with the SoTA planner from the literature [38]. Figure 5 shows the experimental setup and the target poses on the strain map, together with the optimal trajectories that were delivered to our test subject by the robot in both cases. For both executions, the robot started at the shoulder pose marked by the blue dot in the low-strain zone. First, it moved the patient to the pose marked by the orange dot and then toward the state indicated by the green dot to increase the level of strain in a controlled manner and augment the subject’s range of motion. Finally, the robot moved the subject back into the low-strain zone denoted by the pink dot.
BATON’s resulting trajectory avoided higher strain regions (Fig. 5, top left), navigating the rather large range of motion with lower strain overall. On the other hand, the SoTA method produced trajectories that tend to cross zones of higher strain, and lead to oscillations in the subject’s movement (especially evident towards the last pose). The bottom row of Fig. 5 (left) shows the smoothness of the resulting movements in terms of human acceleration magnitude, with the SOTA approach producing consistently higher accelerations with respect to BATON. The evolution of the interaction forces between our subject and the robot is also shown for both approaches (bottom right), with BATON’s trajectory requiring lower interaction forces, especially during the last part of the movement.
V-C Online adaptations (active subject)
@@
Figure 6: The simulated effects of infraspinatus muscle activation ($\alpha_{\text{is}}$) on the optimal trajectory with the same target pose, when $\alpha_{\text{is}}$ follows different activation profiles (green: decreasing ramp, white: step) or remains constant (blue). Large activation steps are used for ease of visualization, and all of the resulting trajectories are projected on the strain map corresponding to the highest $\alpha_{\text{is}}$ for comparison to the right. Our musculoskeletal trajectory optimization is capable of accounting for sudden variations in the strain level, as a consequence, e.g., of varying muscle activation.
V-C1 Simulations
@@
First, we investigated the responsiveness and adaptability of our control scheme when the strain maps change as a consequence of artificial changes in infraspinatus activations. Owing to the shorter time horizon over which we ran our trajectory optimization, BATON was able to achieve a trajectory re-planning with an update rate of roughly 10 Hz. This permits us to generate new trajectories that can account for strain map variations, allowing us to control human movement in a closed-loop manner.
@@
The effect of two different muscle activation changes is visible in Fig. 6: a sudden step increase in activation (white trajectory) and a gradual decrease in activation (green trajectory). These two cases are shown alongside the two boundary cases where activation is kept constantly high or low (resulting in the blue trajectories in the figure). We visualize all of the trajectories jointly on a single strain map, representing the strain map defined by the highest activation, set at $\alpha_{\text{is}}=0.25$ here (Fig. 6, right). Re-projecting all of the trajectories on the same plane highlights differences due to the gradual and sudden variations of the activation level, bounded by the two cases in which the activation was kept constant.
@@
Figure 7: Real-time trajectory optimization allows to account for sudden variations in the strain level. We show here how the optimal trajectory changes as a consequence of rapid variations of (fabricated) strain maps. The trajectory optimization starts on a completely safe strain map, generating a motion directly toward the goal (left). Sudden variations of the maps cause the reference to avoid higher-strain zones (center and right), generating a trajectory that deviates from the shortest path adapting to the new scenarios, and prioritizing safety.
@@
To understand the effects of more sudden and large changes in activation, for example, due to a strong reflex response, we tested cases in which BATON is presented with larger, more localized, and sudden variations in the strain maps (see Fig. 7). First, the strain map on which to plan has low strains, so the optimal trajectory points directly to the goal (Fig. 7, left). We then introduced a sudden large local variation in the strain map topography, so that higher strain appears along the direction of movement. BATON reacted to the variation by re-planning a new trajectory that keeps evolving towards the lower-strain region, deviating from the older plan proposed at the previous time step (white dots in the figure). Similarly, we introduced a second sudden large local variation in the strain map, observing a new trajectory adaptation that allows the simulated movement to remain safe.
V-C2 Real system
Having explored and compared the real-time adaptation of the reference optimal rehabilitation trajectories in simulation, we proceeded to demonstrate this in a real robotic-assisted experiment. Here, online estimates for all of the muscles in the human model were achieved through the RMR solver (Sec. III-A) at 20 Hz. Based on the estimated infraspinatus activation $\hat{\alpha}_{\text{is}}$, the strain map considered by trajectory optimization was dynamically updated at runtime. The subject followed the robot-led movement twice, between the same starting and goal shoulder poses. In the first condition, they remained passive as the robot moved their arm, and the strain map topography was mostly static during execution (Fig. 8, first row). In the second condition instead, they were instructed to perform an external rotation of their arm, resisted by the robot, to elicit infraspinatus activation (Fig. 8, second row). Selected snapshots of the two movements, visualized on the time-varying strain maps, reveal BATON’s adaptations to muscle activation changes. In particular, the two executions lead to a difference of about $10^{\circ}$ in human poses, as a result of adaptations of the human movements to traverse lower tendon strain areas defined by the time-dependent muscle activity values (Fig. 8, third row) resulting from increased human-robot contact wrenches (Fig. 8, bottom row).
V-D Computational performances